Calculate the pOH of a solution that results from mixing 13 mL of 0.12 M methylamine (CH3NH2) with 14 mL of 0.1 M CH3NH3Cl. The Kb value for CH3NH2 is 3.6 x 10-4.
Answer #1

Concentration after mixing = mol of component / (total volume)

M(CH3NH3+) after mixing = M(CH3NH3+)*V(CH3NH3+)/(total volume)

M(CH3NH3+) after mixing = 0.12 M*13.0 mL/(13.0+14.0)mL

M(CH3NH3+) after mixing = 5.778*10^-2 M

Concentration after mixing = mol of component / (total volume)

M(CH3NH2) after mixing = M(CH3NH2)*V(CH3NH2)/(total volume)

M(CH3NH2) after mixing = 0.1 M*14.0 mL/(14.0+13.0)mL

M(CH3NH2) after mixing = 5.185*10^-2 M

Kb = 3.6*10^-4

pKb = - log (Kb)

= - log(3.6*10^-4)

= 3.444

use:

pOH = pKb + log {[conjugate acid]/[base]}

= 3.444+ log {5.185*10^-2/5.778*10^-2}

= 3.397

Answer: 3.40